OpenStudy (cheesecakekitten):

NONE OF YOU GET A MEDAL LOLOLOL I'M GOING TO BRAINLY BECauz it's SOOOOO MUCH BETTTER LOLLOLLOL

OpenStudy (cheesecakekitten):

SRSLY THO I NEED HALP

OpenStudy (cheesecakekitten):

Use the method of completing the square to transform the quadratic equation into the equation form (x – p)^2 = q. 12 - 8x^2 + x^4 = 0 A) (x2 - 4)2 = -4 B) (x2 - 4)2 = 4 C) (x2 - 2)2 = -4 D) (x2 - 2)2 = 4

OpenStudy (osgood):

Lol

OpenStudy (anonymous):

\[x^4-8x^2+12+(-8/2)^2-(-8/2)^2=0\] \[x^4-8x^2+16=4\] \[\left( x^2-4 \right)^2=4\]

OpenStudy (mathstudent55):

\(12 - 8x^2 + x^4 = 0\) \(x^4 - 8x^2 + 12 = 0\) \(x^4 - 8x^2 = -12\) \(x^4 - 8x^2 + \left( \dfrac{-8}{2} \right)^2= -12 + \left( \dfrac{-8}{2} \right)^2\) \(x^4 - 8x^2 + ( -4)^2= -12 + ( -4)^2\) \(x^4 - 8x^2 + 16 = -12 + 16\) \((x^2 - 4)^2 = 4\)